OpenStudy (anonymous):

What is the 25th term of the arithmetic series 9, 15, 21, ...? (3 points) 151 152 153 154

hartnn (hartnn):

do u know the n'th term formula for arithmetic series?

OpenStudy (anonymous):

I know the formula i just always have a hard time plugging numbers in

hartnn (hartnn):

ok, 1st term a1 = 9 common difference d=15-9 = 6 n=25. plug them in and tell me what u get ?

OpenStudy (anonymous):

but you already put what they equal??

hartnn (hartnn):

an = a+(n-1)d so, a25 = 9+(25-1)6 can u find a25 ?>

OpenStudy (anonymous):

33*6 then?

OpenStudy (anonymous):

but that isn't one of the answers I am doing something wrong

hartnn (hartnn):

9+24*6=...?

OpenStudy (anonymous):

you multiplt 24 by 6 first then add 9

hartnn (hartnn):

first you multiply.

OpenStudy (anonymous):

oh right. PEMDAS

OpenStudy (anonymous):

153

OpenStudy (anonymous):

very good!

hartnn (hartnn):

yes.